OpenStudy (supernova_sonntag):

How to find range? f(x)=x^2/1-x^2

OpenStudy (supernova_sonntag):

I know the answer but I need to understand how to do it.

OpenStudy (prathamesh_m):

There's actually no fixed method or algorithm for finding the range. You usually do it using the domain, so first find the domain.

OpenStudy (supernova_sonntag):

Domain: \[(-\infty,-1) \cup(-1,1)\cup(1,\infty)\] \[x \neq1,-1\]

OpenStudy (supernova_sonntag):

what now?

OpenStudy (anthonyym):

Yes that's the right domain. You can check to see if there's horizontal asymptotes. Check out this link: http://www.coolmath.com/precalculus-review-calculus-intro/precalculus-algebra/18-rational-functions-finding-horizontal-slant-asymptotes-01

OpenStudy (supernova_sonntag):

The horizontal asymptote is -1?

OpenStudy (anthonyym):

Yes, it's y = -1

OpenStudy (anthonyym):

So all y-values will be defined for the graph except when y = -1.

OpenStudy (agent0smith):

Looking at a graph is always a good idea to help with finding range: https://www.google.com/search?q=x%5E2%2F1-x%5E2&oq=x%5E2%2F1-x%5E2&aqs=chrome..69i57&sourceid=chrome&ie=UTF-8#q=x%5E2%2F(1-x%5E2) (you can zoom in or out on google too)

OpenStudy (supernova_sonntag):

Oh, so how would I write the answer? Like \[(-\infty,-1) \cup[0,\infty)\]

OpenStudy (agent0smith):

From the graph, that looks exactly right

OpenStudy (supernova_sonntag):

okay cool. However, my text book answer page says that it is \[(-\infty,\infty)\cup[0,\infty)\]

OpenStudy (supernova_sonntag):

is that the same thing?

OpenStudy (agent0smith):

That makes absolutely no sense as a range, haha the first part says the range is all real numbers, -infinity to infinity

OpenStudy (agent0smith):

I'd bet it was a mistake and the infinity should be -1

OpenStudy (supernova_sonntag):

I agree. Thank you!